# Finely Stratified Rerandomization Designs

### v2

@@ 6 Restoring Normality
In this section, we study optimal linearly adjusted GMM estimation under stratified rerandomization. We show that, to first order, optimal ex-post linear adjustment completely removes the impact of the acceptance region $A$ and imbalance term $R_{A}$, restoring asymptotic normality. This enables standard t-statistic and Wald-test based inference on the parameters $\theta_{n}$ and $\theta_{0}$, provided in Section 7 below.
@@
Suppose $D_{1:n}$ as in Definition 2.1 with $A=\mathbb{R}^{d_{h}}$. Require Assumption 3.2. Then we have $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V(\alpha))$ with $V(\alpha)=v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\alpha^{\prime}w|\psi)]$ and $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{0})\Rightarrow\mathcal{N}(0,V_{\phi}+V(\alpha))$.
A version of this result was given in Cytrynbaum, (2023) for the special case $\theta_{0}=\operatorname{ATE}$. Motivated by Proposition 6.2, we define the optimal linear adjustment coefficient as the minimizer of the asymptotic variance $V(\alpha)$, in the positive semidefinite sense.
@@
$\alpha_{0}\in\operatorname*{argmin}_{\alpha\in\mathbb{R}^{d_{w}\times d_{g}}}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\alpha^{\prime}w|\psi)].$ (6.1)
Note that if $w=h$ then $\alpha_{0}=\gamma_{0}$, as in Theorem 3.5. If $E[\operatorname{Var}(w|\psi)]\succ 0$, then the unique minimizer of Equation 6.1 is the partially linear regression coefficient $\alpha_{0}=E[\operatorname{Var}(w|\psi)]^{-1}E[\operatorname{Cov}(w,\Pi a(W,\theta_{0})|\psi)]$. Observe that the optimal adjustment coefficient $\alpha_{0}$ varies with the stratification variables $\psi$, as observed in Cytrynbaum, (2024) and Bai et al., (2023) for ATE estimation. The main result of this section shows that adjustment by a consistent estimate of $\alpha_{0}$ restores asymptotic normality.
@@
Two-step Adjustment. For nonlinear models, the coefficient $\alpha_{0}$ may depend on the unknown parameter $\theta_{0}$. This suggests a two-step adjustment strategy, where we
@@
Suppose $D_{1:n}$ is as in Definition 2.1. Require Assumption 3.1, 3.2. Assume that $E[\operatorname{Var}(w|\psi)]\succ 0$. Define $\widehat{\alpha}=v_{D}E_{n}[\check{w}_{i}\check{w}_{i}^{\prime}]^{-1}E_{n}[H_{i}\check{w}_{i}\widehat{g}_{i}^{\prime}]\widehat{\Pi}^{\prime}$. Then $\widehat{\alpha}=\alpha_{0}+o_{p}(1)$.
In some cases, $\alpha_{0}$ is independent of $\theta_{0}$. For example, if $a(W,\theta)=a_{1}(\psi,\theta)+a_{2}(W)$ then $\alpha_{0}=E[\operatorname{Var}(w|\psi)]^{-1}E[\operatorname{Cov}(w,\Pi a_{2}(W)|\psi)]$ does not depend on $\theta_{0}$. In such cases, one-step optimal adjustment is possible.
@@
One-step adjustment is possible in many linear GMM problems, including the best linear predictor of treatment effects parameter in Example 3.11.
Example 6.6 (Treatment Effect Heterogeneity).
Continuing Example 3.11 with score $g(Y,D,X,\theta)=(HY-X^{\prime}\theta)X$ and $\theta_{n}=\operatorname*{argmin}_{\theta}E_{n}[(\tau_{i}-X_{i}^{\prime}\theta)^{2}]$, recall that $a(W,\theta_{0})=\bar{Y}X$ and $\Pi=E[XX^{\prime}]^{-1}$. Letting $\theta=0$ gives $g(Y,D,X,0)=HYX$. By Corollary 6.5, we have $\widehat{\alpha}=\alpha_{0}+o_{p}(1)$ for the adjustment coefficient matrix
$\widehat{\alpha}=v_{D}E_{n}[\check{w}_{i}\check{w}_{i}^{\prime}]^{-1}E_{n}[H_{i}^{2}Y_{i}\check{w}_{i}X_{i}^{\prime}]E_{n}[X_{i}X_{i}^{\prime}]^{-1}.$ (6.2)
This allows us to estimate treatment effect heterogeneity relative to a low-dimensional vector of important covariates $X$, adjusting optimally for a larger set of covariates $w$ ex-post in order to improve precision, as well as restore asymptotic normality when $A\not=\mathbb{R}^{d_{h}}$. In the case $X=1$ (SATE estimation), $\widehat{\alpha}$ is equivalent to the “tyranny-of-the-minority” style estimator proposed in Cytrynbaum, (2023).
7 Inference
@@
In this section, we provide novel methods for inference on general causal parameters under stratified rerandomization designs. We make crucial use of asymptotic normality of the optimally adjusted estimator $\widehat{\theta}_{adj}$, shown in Theorem 6.3. For the superpopulation parameter $\theta_{0}$, we provide asymptotically exact inference methods. The asymptotic variance for estimating the finite population parameter $\theta_{n}$ is generally not identified. In this case, we provide conservative variance estimation that still reflects the precision gains due to stratification and rerandomization.
7.1 Asymptotically Exact Inference
To define our variance estimator, we begin with some definitions. Let $\mathcal{S}_{n}$ denote the set of groups constructed in Definition 2.1. For each $s\in\mathcal{S}_{n}$ define the centroid $\bar{\psi}_{s}=|s|^{-1}\sum_{i\in s}\psi_{i}$. Let $\nu:\mathcal{S}_{n}\to\mathcal{S}_{n}$ be a bijective matching between groups satisfying $\nu(s)\not=s$, $\nu^{2}=\operatorname{Id}$, and the homogeneity condition
$\frac{1}{n}\sum_{s\in\mathcal{S}_{n}}|\bar{\psi}_{s}-\bar{\psi}_{\nu(s)}|_{2}^{2}=o_{p}(1).$ (7.1)

### v3

@@ 6 Restoring Normality
In this section, we study optimal linearly adjusted GMM estimation under stratified rerandomization. We show that, to first order, optimal ex-post linear adjustment completely removes the impact of the acceptance region $A$ and imbalance term $R_{A}$, restoring asymptotic normality. This enables standard t-statistic and Wald-test based inference on the parameters $\theta_{n}$ and $\theta_{0}$ under stratified rerandomization designs, provided in Section 7 below. We also describe a novel form of double robustness to covariate imbalances from combining rerandomization with ex-post adjustment.
@@
Suppose $D_{1:n}$ as in Definition 2.1 with $A=\mathbb{R}^{d_{h}}$. Require Assumption 3.2. Then we have $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a}(\alpha))$ with $V_{a}(\alpha)=v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\alpha^{\prime}w|\psi)]$ and $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{0})\Rightarrow\mathcal{N}(0,V_{\phi}+V_{a}(\alpha))$.
A version of this result was given in Cytrynbaum, 2024a for the special case $\theta_{0}=\operatorname{ATE}$. Motivated by Proposition 6.2, we define the optimal linear adjustment coefficient as the minimizer of the asymptotic variance $V_{a}(\alpha)$, in the positive semidefinite sense.
@@
$\alpha_{0}\in\operatorname*{argmin}_{\alpha\in\mathbb{R}^{d_{w}\times d_{\theta}}}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\alpha^{\prime}w|\psi)].$ (6.1)
Note that if $w=h$ then $\alpha_{0}=\gamma_{0}$ in Theorem 3.5. If $E[\operatorname{Var}(w|\psi)]\succ 0$, then the unique minimizer of Equation 6.1 is the partially linear regression coefficient matrix $\alpha_{0}=E[\operatorname{Var}(w|\psi)]^{-1}E[\operatorname{Cov}(w,\Pi a(W,\theta_{0})|\psi)]$. Observe that $\alpha_{0}$ varies with the stratification variables $\psi$, as observed in Cytrynbaum, 2024b and Bai et al., 2024a for the case of ATE estimation. The main result of this section shows that adjustment by a consistent estimate of $\alpha_{0}$ restores asymptotic normality.
@@
Two-step Adjustment. For nonlinear models, the optimal coefficient $\alpha_{0}$ may depend on the unknown parameter $\theta_{0}$. This suggests a two-step adjustment strategy:
@@
$\widehat{\alpha}=v_{D}E_{n}[\check{w}_{i}\check{w}_{i}^{\prime}]^{-1}\left[\operatorname{Cov}_{n}(\check{w}_{i},\widehat{\Pi}\widehat{g}_{i}|D_{i}=1)-\operatorname{Cov}_{n}(\check{w}_{i},\widehat{\Pi}\widehat{g}_{i}|D_{i}=0)\right].$ (6.2)
@@
Suppose $D_{1:n}$ is as in Definition 2.1. Require Assumption 3.1, 3.2. Assume that $E[\operatorname{Var}(w|\psi)]\succ 0$. Then $\widehat{\alpha}=\alpha_{0}+o_{p}(1)$.
In some cases, $\alpha_{0}$ may not depend on $\theta_{0}$. For example, if $a(W,\theta)=a_{1}(\psi,\theta)+a_{2}(W)$ then $\alpha_{0}=E[\operatorname{Var}(w|\psi)]^{-1}E[\operatorname{Cov}(w,\Pi a_{2}(W)|\psi)]$. In such cases, one-step optimal adjustment is possible.
@@
One-step adjustment is possible in many linear GMM problems, including the best linear predictor of treatment effect heterogeneity parameter in Example 3.10.
Example 6.6 (Adjusting CATE Estimate).
Continuing Example 3.10, suppose we want to estimate treatment effect heterogeneity relative to an important covariate $X$, while adjusting optimally for larger set of measured covariates $w$ to both improve precision and restore asymptotic normality under rerandomization. For GMM score $g(Y,D,X,\theta)=(HY-X^{\prime}\theta)X$ we have $\theta_{n}=\operatorname*{argmin}_{\theta}E_{n}[(Y_{i}(1)-Y_{i}(0)-X_{i}^{\prime}\theta)^{2}]$. Then $a(W,\theta)=\bar{Y}X$ and $\Pi=E[XX^{\prime}]^{-1}$. Letting $\theta=0$ gives $g(Y,D,X,0)=HYX$. After some algebra, Corollary 6.5 shows that $\widehat{\alpha}=\alpha_{0}+o_{p}(1)$ for adjustment coefficient
$\widehat{\alpha}$ $=E_{n}[\check{w}_{i}\check{w}_{i}^{\prime}]^{-1}\left[(1-p)\operatorname{Cov}_{n}(\check{w}_{i},Y_{i}X_{i}|D_{i}=1)+p\operatorname{Cov}_{n}(\check{w}_{i},Y_{i}X_{i}|D_{i}=0)\right]E_{n}[X_{i}X_{i}^{\prime}]^{-1}.$
We apply this adjustment in our empirical application to estimating treatment effect heterogeneity for the experiment in Angrist et al., (2013) in Section 9 below.
6.1 Double Robustness from Rerandomization
Refer to caption Refer to caption
Figure 3: MSE for Adjustment vs. Rerandomization, $\theta_{n}=\operatorname{SATE}$ and $n\in\{150,500\}$.
Theorem 6.3 shows that stratified rerandomization has (approximately) the same first-order efficiency as optimal ex-post linear adjustment tailored to both the stratification and GMM problem. [^30] For the case without stratification, this equivalence was originally shown in Li et al., (2018). However, our simulations and empirical application show that in finite samples stratified rerandomization can perform significantly better than ex-post adjustment, and further efficiency gains are possible by combining both methods. In this subsection, we provide a brief theoretical justification for this phenomenon, showing that combining rerandomization and adjustment provides a novel form of double robustness to covariate imbalances.
For simplicity, consider the case of $\operatorname{SATE}$ estimation with $\psi=1$. For the difference of means estimator $\widehat{\theta}$, by a simple calculation $\widehat{\theta}-\operatorname{SATE}=E_{n}[\bar{Y}_{i}|D_{i}=1]-E_{n}[\bar{Y}_{i}|D_{i}=0]$. Let $\bar{Y}=c+\gamma_{0}^{\prime}h+e$ with $e\perp(1,h)$ be the decomposition from Equation 5.1. Then we can decompose the estimation error $\widehat{\theta}-\operatorname{SATE}$ into imbalances in $h$ and $e$:
$\sqrt{n}(\widehat{\theta}-\operatorname{SATE})=\sqrt{n}\gamma_{0}^{\prime}(\bar{h}_{1}-\bar{h}_{0})+\sqrt{n}(\bar{e}_{1}-\bar{e}_{0}).$ (6.3)
Rerandomizing until $\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$ is small shrinks the first imbalance term, ideally making its variance negligible relative to $\operatorname{Var}(e)=\operatorname{Var}(\bar{Y}-\gamma_{0}^{\prime}h)$. Suppse $h=w$ so the adjustment coefficient $\alpha_{0}=\gamma_{0}$. Writing $\widehat{\alpha}=\widehat{\gamma}$, we have $\widehat{\theta}_{adj}=\widehat{\theta}-\widehat{\gamma}^{\prime}E_{n}[H_{i}h_{i}]=\widehat{\theta}-\widehat{\gamma}^{\prime}(\bar{h}_{1}-\bar{h}_{0})$,
$\sqrt{n}(\widehat{\theta}_{adj}-\operatorname{SATE})=\sqrt{n}(\gamma_{0}-\widehat{\gamma})^{\prime}(\bar{h}_{1}-\bar{h}_{0})+\sqrt{n}(\bar{e}_{1}-\bar{e}_{0}).$ (6.4)
This decomposition shows a novel form of double robustness from combining rerandomization with ex-post adjustment. If the estimation error $\gamma_{0}-\widehat{\gamma}$ is large, then the first imbalance term above may still be negligible as long as we rerandomized until $\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$ is small enough. For example, in the LHS of Figure 3 the coefficient $\gamma_{0}$ is not estimated well for small $n$ and large $\dim(w)$, so adjustment without rerandomization performs poorly. This effect is exacerbated by stratification, since the partialling operation $\check{w}_{i}=w_{i}-\sum_{j\in s(i)}w_{j}$ tends to decrease the variance of the regressors $w_{i}$, making estimation of $\alpha_{0}$ more difficult. [^31] For example, in our empirical application the condition number of the design matrix $E_{n}[\check{w}_{i}\check{w}_{i}]$ increases as we stratify more finely. However, combining adjustment and rerandomization still performs well due to double robustness.
Equation 6.4 has a “product of errors” structure, similar to the product of nuisance estimation errors for doubly-robust estimators in the literature on Neyman orthogonal estimating equations (e.g. Chernozhukov et al., (2017)). This shows that even when both $\widehat{\gamma}-\gamma_{0}$ and $\bar{h}_{1}-\bar{h}_{0}$ are small, we can get an extra benefit from combining the two methods. This is shown in the RHS of Figure 3 with $n=500$, where adjustment is competitive, but rerandomization still performs better, and stratification + rerandomization is even more efficient due to this product structure. Such double robustness also holds for more general casual parameters and GMM estimators. Let $h=w$ and consider the partially linear decomposition $\Pi a(W,\theta_{0})=\gamma_{0}^{\prime}h+t(\psi)+e$ with $e\perp h$ and $E[e|\psi]=0$ and $\bar{t}_{d}=E_{n}[t(\psi_{i})|D_{i}=d]$. Our work shows that
$\widehat{\theta}_{adj}-\theta_{n}$ $=(\gamma_{0}-\widehat{\gamma})^{\prime}(\bar{h}_{1}-\bar{h}_{0})+(\bar{t}_{1}-\bar{t}_{0})+(\bar{e}_{1}-\bar{e}_{0})+o_{p}(n^{-1/2})$
$=(\gamma_{0}-\widehat{\gamma})^{\prime}(\bar{h}_{1}-\bar{h}_{0})+(\bar{e}_{1}-\bar{e}_{0})+o_{p}(n^{-1/2}).$
The second equality is a consequence of fine stratification on $\psi$.
Summarizing, this discussion highlights a double robustness property that explains the additional finite-sample precision gains from combining rerandomization with ex-post adjustment. This effect is likely to be especially important in regimes where the optimal adjustment coefficient $\alpha_{0}$ is poorly estimated, such as for small $n$, large $\dim(w)$, ill-conditioned design matrix $E_{n}[\check{w}_{i}\check{w}_{i}^{\prime}]\approx E[\operatorname{Var}(w|\psi)]$. A full theory of high-dimensional stratification, rerandomization, and ex-post adjustment is beyond the scope of the current work, but this is an interesting area for future research. [^32] There are analytical complications from conditioning on $\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})\in A$ e.g. with $\dim(h)$ growing. A recent breakthrough on this question was achieved by the careful analysis of Wang and Li, (2022) for the case of complete rererandomization.
7 Variance Bounds and Inference Methods
@@
In this section, we provide methods for inference on generic causal parameters under stratified rerandomization designs. We make crucial use of asymptotic normality of the optimally adjusted GMM estimator $\widehat{\theta}_{adj}$ developed in the previous section. The asymptotic variance for estimating the finite population parameter $\theta_{n}$ is generally not identified. To enable inference, we provide novel identified upper bounds on the variance, allowing for conservative inference that still reflects the precision gains from stratified rerandomization. The asymptotic variance for estimating the superpopulation parameter $\theta_{0}$ is identified, and in this case we provide asymptotically exact inference methods.
7.1 Variance Bounds
First, we briefly review the classical variance bounds for $\theta_{n}=\operatorname{SATE}$ estimation under completely randomized assignment. In this case, we have $\sqrt{n}(\widehat{\theta}-\operatorname{SATE})\Rightarrow\mathcal{N}(0,V_{a})$ with $V_{a}=\operatorname{Var}(D)^{-1}\operatorname{Var}(\bar{Y})$ for $\bar{Y}=(1-p)Y(1)+pY(0)$. The variance $\operatorname{Var}(\bar{Y})\propto\operatorname{Cov}(Y(1),Y(0))$. Since $Y(1)$ and $Y(0)$ are never simultaneously observed, $V_{a}$ is not identified. Let $\sigma^{2}_{d}=\operatorname{Var}(Y(d))$ and $\tau=Y(1)-Y(0)$. The Cauchy-Schwarz inequality $|\operatorname{Cov}(Y(1),Y(0))|\leq\sigma_{1}\sigma_{0}$ and some algebra produces the bounds
$V_{a}=\frac{\sigma^{2}_{1}}{p}+\frac{\sigma^{2}_{0}}{1-p}-\operatorname{Var}(\tau)\leq\frac{\sigma^{2}_{1}}{p}+\frac{\sigma^{2}_{0}}{1-p}-(\sigma_{1}-\sigma_{0})^{2}\leq\frac{\sigma^{2}_{1}}{p}+\frac{\sigma^{2}_{0}}{1-p}.$ (7.1)
Both upper bounds were proposed in Neyman, (1990). Theorem 7.1 below extends the sharper bound to generic finite population causal parameters, accounting for both design-time stratified rerandomization and optimal ex-post adjustment.
To develop the bounds, recall from Theorem 6.3 that $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{n})\Rightarrow N(0,V_{a}^{adj})$ with $V_{a}^{adj}=v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\alpha_{0}^{\prime}w|\psi)]$, where $\alpha_{0}=E[\operatorname{Var}(w|\psi)]^{-1}E[\operatorname{Cov}(w,\Pi a(W,\theta_{0})|\psi)]$ was the optimal adjustment coefficient. By definition, $\Pi a(W,\theta_{0})=v_{D}\Pi(g_{1}(W,\theta_{0})-g_{0}(W,\theta_{0}))$. Then the adjustment coefficient may be expanded as $\alpha_{0}=\beta_{1}-\beta_{0}$ for coefficients $\beta_{d}=E[\operatorname{Var}(w|\psi)]^{-1}E[\operatorname{Cov}(w,v_{D}\Pi g_{d}(W,\theta_{0})|\psi)]$. Denote $g_{d}=g_{d}(W,\theta_{0})$ and define the “within-arm” influence functions $m_{d}\equiv v_{D}\Pi g_{d}-\beta_{d}^{\prime}w$. Tighter bounds are possible by targeting a fixed scalar contrast $c^{\prime}\theta_{n}$ for some $c\in\mathbb{R}^{d_{\theta}}$. From Theorem 6.3, we have $\sqrt{n}(c^{\prime}\widehat{\theta}_{adj}-c^{\prime}\theta_{0})\Rightarrow N(0,V_{a}^{adj}(c))$ for $V_{a}^{adj}(c)=c^{\prime}V_{a}^{adj}c$. In terms of $m_{d}$, this is
$V_{a}^{adj}(c)$ $=c^{\prime}v_{D}^{-1}E[\operatorname{Var}(v_{D}\Pi(g_{1}-g_{0})-(\beta_{1}-\beta_{0})^{\prime}w|\psi)]c$
$=v_{D}^{-1}E[\operatorname{Var}(c^{\prime}m_{1}-c^{\prime}m_{0}|\psi)].$
Similarly to above, $V_{a}^{adj}(c)\propto E[\operatorname{Cov}(c^{\prime}m_{1},c^{\prime}m_{0}|\psi)]$ where the cross-term is generically not identified, since $m_{1}$ and $m_{0}$ are not simultaneously observed. However, denoting $\tilde{\sigma}^{2}_{d}(c)=E[\operatorname{Var}(c^{\prime}m_{d}|\psi)]$ we have the following simple upper bound:
Theorem 7.1 (Variance Bounds).
Under the conditions of Theorem 6.3, we have
$V_{a}^{adj}(c)$ $\leq v_{D}^{-1}(\tilde{\sigma}_{1}(c)+\tilde{\sigma}_{0}(c))^{2}=v_{D}^{-1}\left(\frac{\tilde{\sigma}^{2}_{1}(c)}{1-p}+\frac{\tilde{\sigma}^{2}_{0}(c)}{p}\right)-\left(\frac{\tilde{\sigma}_{1}(c)}{1-p}-\frac{\tilde{\sigma}_{0}(c)}{p}\right)^{2}.$
We provide a consistent estimator of the bound $\bar{V}_{a}^{adj}(c)$ in Section 7.2 below. The next example shows how Theorem 7.1 generalizes the classical Neyman bounds for the simple case of inference on $\theta_{n}=\operatorname{SATE}$ under pure stratified randomization ($A=\mathbb{R}^{d_{h}}$) and optimal ex-post adjustment.
Example 7.2 (Pure Stratification).
Let $\theta_{n}=\operatorname{SATE}$ so $c=1$. Then for $H=\frac{D-p}{p-p^{2}}$ and GMM score $g(D,Y,\theta)=HY-\theta$ have $\Pi=1$ and $v_{D}\Pi g_{1}=(p-p^{2})Y(1)/p=(1-p)Y(1)$. Then $\beta_{1}=(1-p)\delta_{1}$ for $\delta_{1}=\operatorname*{argmin}_{\delta}E[\operatorname{Var}(Y(1)-\delta^{\prime}w|\psi)]$ we have $m_{1}=(1-p)(Y(1)-\delta_{1}^{\prime}w)$. Similarly, $m_{0}=p(Y(0)-\delta_{0}^{\prime}w)$ with $\delta_{0}=\operatorname*{argmin}_{\delta}E[\operatorname{Var}(Y(0)-\delta^{\prime}w|\psi)]$. Plugging into the second expression in Theorem 7.1, the optimally adjusted variance $V_{a}^{adj}=\min_{\gamma}v_{D}^{-1}E[\operatorname{Var}(\bar{Y}-\gamma^{\prime}w|\psi)]$ is bounded above by
$\bar{V}_{a}^{adj}$ $=\frac{E[\operatorname{Var}(Y(1)-\delta_{1}^{\prime}w|\psi)]}{p}+\frac{E[\operatorname{Var}(Y(0)-\delta_{0}^{\prime}w|\psi)]}{1-p}$
$-(E[\operatorname{Var}(Y(1)-\delta_{1}^{\prime}w|\psi)]^{1/2}-E[\operatorname{Var}(Y(0)-\delta_{0}^{\prime}w|\psi)]^{1/2})^{2}.$
For unadjusted complete randomization ($\psi=1$, $w=0$), we recover the sharper Neyman bound in Equation 7.1. If $\psi\not=1$ and $w=0$, we get a novel “finely stratified” bound:
$\bar{V}_{a}=\frac{E[\operatorname{Var}(Y(1)|\psi)]}{p}+\frac{E[\operatorname{Var}(Y(0)|\psi)]}{1-p}-(E[\operatorname{Var}(Y(1)|\psi)]^{1/2}-E[\operatorname{Var}(Y(0)|\psi)]^{1/2})^{2}.$
Remark 7.3 (Covariate-Assisted Bounds by Design).
In some contexts, it is possible to use covariate information to tighten finite population variance bounds, e.g. as in Abadie et al., (2020). For example, under complete randomization with $\theta_{n}=\operatorname{SATE}$, the non-identified $\operatorname{Cov}(Y(1),Y(0))=E[\operatorname{Cov}(Y(1),Y(0)|\psi)]+\operatorname{Cov}(E[Y(1)|\psi],E[Y(0)|\psi])\equiv v_{1}+v_{2}$ by law of total covariance. Only $v_{1}$ is non-identified, while $v_{2}$ can be consistently estimated using $\psi$. In our context, however, the term $v_{2}$ is already removed from the asymptotic variance due to stratified randomization of $D_{1:n}$. More generally, under stratified rerandomization with adjustment, $V_{a}\propto v_{1}=E[\operatorname{Cov}(Y(1)-\delta_{1}^{\prime}w,Y(0)-\delta_{0}^{\prime}w|\psi)]$, so covariate-assisted tightening happens “automatically” by design. Relative to the papers above, our work provides a tighter upper bound on $v_{1}$ even after covariate-assistance, corresponding to the sharper Neyman bound in Equation 7.1.
Remark 7.4 (Sharp Bounds).
For $\theta_{n}=\operatorname{SATE}$ estimation under completely randomized assignment, Aronow et al., (2014) derive sharp upper bounds on the variance $V_{a}=v_{D}^{-1}\operatorname{Var}(\bar{Y})$. In principle, such bounds could be extended to the more general designs and estimators in our current setting. However, this construction and the associated variance estimators are quite involved, so we leave this significant extension to future work. [^33] Alternatively, note $E[\operatorname{Cov}(Y(1),Y(0)|\psi)]\leq E[\sigma_{1}(\psi)\sigma_{0}(\psi)]\leq E[\sigma^{2}_{1}(\psi)]^{1/2}E[\sigma^{2}_{0}(\psi)]^{1/2}$. Theorem 7.1 uses the second bound, which we prefer since it can be naturally estimated using the stratification. The first bound could be tighter for large heteroskedasticity, but requires additional nonparametric estimation.
7.2 Inference on the Finite Population Parameter
Building on the previous section, we construct a consistent estimator of the variance upper bound $\bar{V}_{a}^{adj}(c)$, enabling asymptotically conservative inference on linear contrasts of the finite population parameter $c^{\prime}\theta_{n}$ under general designs.